OpenStudy (anonymous):

How do I solve e^2x=x+2

OpenStudy (anonymous):

Think of it as: \[\ln \left( x+2 \right)=2x\]You're probably being asked to play with the properties of logarithms; look it up in whatever reference material you're using, think about it for a minute, and the answer should become clear.